# CooHOI: Learning Cooperative Human-Object Interaction with Manipulated Object Dynamics

### v1

@@ 4.2 Evaluation on Object Carrying Tasks
We evaluate our policy in both single-agent and two-agent scenarios. Table 1 presents the performance statistics for the carrying task. Consistent with the settings described in Interphys Hassan et al. (2023), we randomize the weight and distance as outlined in Section4.1. It is important to note that InterPhys Hassan et al. (2023) includes the weight information of objects as part of the observation, enabling it to achieve a broader weight range. To ensure result comparability, as shown in Line 2, we also incorporated this information into individual training. The results indicate that, with a similar success rate, we increased the transportation distance and improved transportation accuracy, even when using discrete motion data rather than the costly whole-body motion capture data used in InterPhys Hassan et al. (2023). As shown in Line 3, although removing the weight information reduces the weight range the agent can handle, it represents a more realistic setting and provides comparable baseline results for future research. Lines 4-5 show the results in multi-agent scenarios. Without using our CooHOI framework and simply employing parallel training for multi-agent tasks, the training fails. More analysis and visualizations will be provided in the following sections.
Figure 4: Visualization of cooperative carrying in the multi-agent scenario. The second row shows four agents carrying the object. Note that in this scenario, the four agents can only reach this stage and are unable to successfully perform the subsequent put-down operation.
@@
Moreover, we are not satisfied with merely transporting simple boxes. Therefore, we have sampled nearly 40 common everyday objects from Hassan et al. (2021), including three types of single-person objects—Table, Armchair, and HighStool—as well as one type of two-person object, the Sofa, as illustrated in Figure 3 and Line 1 in Figure 4. Additional visualizations of these objects can be found in appendix. Additionally, we present the results of quantitative experiments in Table 2, which indicate that our policy can effectively handle most daily life objects with a high success rate. Due to the greater variability in the shape and weight of sofas, the accuracy in Line 4 shows a slight drop compared to the result in Line 4 of Table 1. It is important to note that the significant variations in object shapes make it challenging to learn a unified object representation with a limited amount of data, especially since our input is merely a simple bounding box. Therefore, for single-person transport, different types of objects need to be trained separately from scratch, similar to how we handle boxes. For multi-person transport, objects like sofas, similar to boxes, require fine-tuning based on the weights obtained from single-person box transport.
@@ 4.3 Ablation Studies
To evaluate and understand the importance of different design choices in CooHOI, we conduct a detailed analysis of scenarios involving single and multiple agents. This includes boundary analysis, which explores factors that lead to a decrease in the success rate during object carrying. To ensure the reliability of our results, our experiments continue to utilize the average of 4 random seeds as previously mentioned.
@@
We conducted upper-limit testing on the multi-agent policy using a similar evaluation method as the single-agent case. As shown in Figure 5, the conclusions for two agents are similar to those for a single agent. The curves in the first figure demonstrate that with an increase in the number of agents, we can easily lift larger and heavier objects that a single person cannot handle, highlighting the necessity of having multiple agents. However, due to the increased complexity of coordinating two agents, boundary conditions have a more pronounced impact on the policy. For instance, objects that are excessively large or small can cause the policy to fail. As shown by the purple circle in the second figure of Figure 5, we conducted a similar experiment to the single-agent scenario. By restricting the object’s width to 1x and scaling the length and height to 1.5x, the success rate increased from 0 to 88.67%. Furthermore, we extended our investigation beyond the two-agent scenario to explore the upper limit of the number of agents by increasing the count to four. As illustrated in the second row of Figure 4, our strategy enables four agents to successfully lift a huge object weighing nearly 100 kg. However, it should be noted that due to the absence of dexterous hands and limited friction, the agents encounter difficulties in stably placing the object down. Addressing this challenge will be a key focus of our future research efforts.
Analysis For CooHOI Framework.
@@
To thoroughly investigate the contribution of CooHOI, we analyzed the results of each method mentioned in Section 3 separately, as shown by the training curves in Figure 6. The first factor is the influence of the Stand Point, which refers to whether an extra point is introduced in front of the object to encourage the agent to walk toward it. During the experiments, we discovered that without this, the agent sometimes fails to approach the shortest edge of the object, resulting in a lower hold reward. This leads to an incomplete lift and the subsequent inability to carry the object. Dynamic Observation is the second factor, indicating whether we use dynamic information as observation for each agent. Without it, the observation for each agent is limited to the state information of the long object. We found that without the dynamics information, the agent just stands in front of the object, seemingly unsure of what to do. Reverse Walk indicates whether the training process includes motion data for walking backward and a reward function focused on learning this movement. We found that if the policy for a single agent is restricted to only forward movement, training with two agents then leads to a deadlock state. As shown in Figure 6, the agents might be able to contact the box, but they cannot carry it to the destination. Initialization refers to whether the two-agent policy is initialized using the single-agent policy and then be fine-tuned. In our experiment, even with extended training duration, training the two-agent policy from scratch still failed to achieve successful carrying, as shown in Figure 6. Based on the results above, the absence of any of the aforementioned methods causes our policy to fall into a locally suboptimal solution, preventing the completion of the transportation task. Moreover, you can also find more interesting visual examples of the above failure cases in the appendix.
@@ 5 Conclusion and Limitation
In this paper, we present Cooperative Human-Object Interaction (CooHOI), a novel framework designed to address cooperative object transporting tasks through a two-phase learning approach. By initially focusing on individual skill mastery via the Adversarial Motion Priors (AMP) method, followed by a strategic transition to multi-agent collaboration using Multi-Agent Proximal Policy Optimization (MAPPO), our approach facilitates a sophisticated interplay of shared dynamics and implicit communication among agents, resulting in an efficient and generalized performance.
However, it is important to note that within the scope of this paper, our agents do not possess active perception capabilities: all object location information is pre-perceived. Additionally, the lack of dexterous hands in our humanoid model limits its performance, preventing the characters from handling objects that are too small or slippery to grip. In future research, we aim to incorporate dexterous hands into our humanoid model to enhance its capabilities and also explore the integration of active perception and navigational abilities.
Supplementary Material
@@ Appendix A Sources and Processing of Motion Data
We collected a total of four types of basic reference motion data, including 9 motions related to walking, 5 related to picking up, 4 related to carrying, and 5 related to putting down. All these data are in SMPL format and recorded at 30 fps over 139 frames. They all originate from the ACCAD subset of the AMASS Mahmood et al. [2019] dataset. Additionally, to ensure the stability of cooperative tasks involving multiple individuals, we included data for sidewalk and reverse carry motions. The sidewalk data comes from the CMU subset within AMASS, while reverse carry data was scarce. Therefore, we created reverse carry data by reversing the process of the carry data. In total, we used 26 motion data as references. Additionally, we performed a simple visualization of the extended objects as in Figure 7, which sampled from dataset Hassan et al. [2021].
@@ B.1 Task Observation
While the root height is measured in the global reference frame, all other components are defined in the frame local to the character. Rotations follow a 6-dimensional representation for continuity Zhou et al. [2019]. The simulated character aligns with Peng et al. [2021, 2022], Hassan et al. [2023], Pan et al. [2023], featuring 12 internally movable joints and a total of 28 degrees of freedom.
@@ B.2 Reward Functions
Follow the formulation of the AMP framework Peng et al. [2021], the style reward $r^{S}$ is calculated according to the discriminator:
@@ Appendix C Implementation Details
Adopting the methodology of AMP Peng et al. [2021], we develop a low-level controller encompassing both policy and discriminator networks. The policy network is bifurcated into a critic and an actor-network, each initiating with a CNN layer and proceeding to two MLP layers configured with [1024, 1024, 512] units. The discriminator network is similarly structured, featuring two MLP layers with [1024, 1024, 512] units. We select PPO Schulman et al. [2017] as the primary reinforcement learning algorithm, coupled with the Adam optimizer Kingma and Ba [2014] at a learning rate of 2e-5. The only difference between the multi-agent setting and the single-agent setting during training is whether a pre-trained weight is loaded. Our experiments are conducted on the IsaacGym simulator Makoviychuk et al. [2021] using a single Nvidia GTX 3090Ti GPU. We run 4096 parallel environments across 15,000 epochs, which takes approximately 15 hours to complete.
@@ C.2 Hyperparameters
Following previous workPeng et al. [2021], Hassan et al. [2023], Pan et al. [2023], we use the Isaac Gym simulator Makoviychuk et al. [2021]. The simulation runs at 60Hz and the control policy runs at 30Hz.
@@
Appendix D Failure case visualization.

### v2

@@ 4.2 Evaluation on Object Carrying Tasks
We evaluate our policy in both single-agent and two-agent scenarios. Table 1 presents the performance statistics for the carrying task. Consistent with the settings described in Interphys [7], we randomize the weight and distance as outlined in Section4.1. It is important to note that InterPhys [7] includes the weight information of objects as part of the observation, enabling it to achieve a broader weight range. To ensure result comparability, as shown in Line 2, we also incorporated this information into individual training. The results indicate that, with a similar success rate, we increased the transportation distance and improved transportation accuracy, even when using discrete motion data rather than the costly whole-body motion capture data used in InterPhys [7]. As shown in Line 3, although removing the weight information reduces the weight range the agent can handle, it represents a more realistic setting and provides comparable baseline results for future research. Lines 4-5 show the results in multi-agent scenarios. Without using our CooHOI framework and simply employing parallel training for multi-agent tasks, the training fails. More analysis and visualizations will be provided in the following sections.
Figure 4: Visualization of cooperative carrying in the multi-agent scenario.
@@
Moreover, we are not satisfied with merely transporting simple boxes. Therefore, we have sampled nearly 40 common everyday objects from [6], including three types of single-person objects—Table, Armchair, and HighStool—as well as one type of two-person object, the Sofa, as illustrated in Figure 3 and Line 1 in Figure 4. Additional visualizations of these objects can be found in appendix. Additionally, we present the results of quantitative experiments in Table 2, which indicate that our policy can effectively handle most daily life objects with a high success rate. Due to the greater variability in the shape and weight of sofas, the accuracy in Line 4 shows a slight drop compared to the result in Line 4 of Table 1. It is important to note that the significant variations in object shapes make it challenging to learn a unified object representation with a limited amount of data, especially since our input is merely a simple bounding box. Therefore, for single-person transport, different types of objects need to be trained separately from scratch, similar to how we handle boxes. For multi-person transport, objects like sofas, similar to boxes, require fine-tuning based on the weights obtained from single-person box transport.
@@ 4.3 Ablation Studies
To evaluate and understand the importance of different design choices in CooHOI, we conduct a detailed analysis of scenarios involving single and multiple agents. This includes boundary analysis, which explores factors that lead to a decrease in the success rate during object carrying. Furthermore, to fully validate the scalability of our method, we also tested and included the results in a four-person scenario. To ensure the reliability of our results, our experiments continue to utilize the average of 4 random seeds as previously mentioned.
@@
We conducted upper-limit testing on the multi-agent policy using a similar evaluation method as the single-agent case. As shown in Figure 5, the conclusions for two agents are similar to those for a single agent. The curves in the first figure demonstrate that with an increase in the number of agents, we can easily lift larger and heavier objects that a single person cannot handle, highlighting the necessity of having multiple agents. However, due to the increased complexity of coordinating two agents, boundary conditions have a more pronounced impact on the policy. For instance, objects that are excessively large or small can cause the policy to fail. As shown by the purple circle in the second figure of Figure 5, we conducted a similar experiment to the single-agent scenario. By restricting the object’s width to 1x and scaling the length and height to 1.5x, the success rate increased from 0 to 88.67%.
Analysis of CooHOI Framework.
@@
To thoroughly investigate the contribution of CooHOI, we analyzed the results of each method mentioned in Section 3 separately, as shown by the training curves in Figure 8. The first factor is the influence of the Stand Point, which refers to whether an extra point is introduced in front of the object to encourage the agent to walk toward it. During the experiments, we discovered that without this, the agent sometimes fails to approach the shortest edge of the object, resulting in a lower hold reward. This leads to an incomplete lift and the subsequent inability to carry the object. Dynamic Observation is the second factor, indicating whether we use dynamic information as observation for each agent. Without it, the observation for each agent is limited to the state information of the long object. We found that without the dynamics information, the agent just stands in front of the object, seemingly unsure of what to do. Reverse Walk indicates whether the training process includes motion data for walking backward and a reward function focused on learning this movement. We found that if the policy for a single agent is restricted to only forward movement, training with two agents then leads to a deadlock state. As shown in Figure 8, the agents might be able to contact the box, but they cannot carry it to the destination. Initialization refers to whether the two-agent policy is initialized using the single-agent policy and then be fine-tuned. In our experiment, even with extended training duration, training the two-agent policy from scratch still failed to achieve successful carrying, as shown in Figure 8. Based on the results above, the absence of any of the aforementioned methods causes our policy to fall into a locally suboptimal solution, preventing the completion of the transportation task. Moreover, you can also find more interesting visual examples of the above failure cases while extending agent numbers from one to two in the appendix.
@@ 5 Conclusion and Limitation
In this paper, we present Cooperative Human-Object Interaction (CooHOI), a framework designed to address cooperative object transporting tasks through a two-phase learning approach. By initially focusing on individual skill mastery via the Adversarial Motion Priors (AMP) method, followed by a strategic transition to multi-agent collaboration using Multi-Agent Proximal Policy Optimization (MAPPO), our approach facilitates a sophisticated interplay of shared dynamics and implicit communication among agents, resulting in an efficient and generalized performance.
However, it is important to note that within the scope of this paper, our huamnoid characters lack dexterous hands, which limit their ability to manipulate slippery objects or perform more precise actions. In future research, we aim to incorporate dexterous hands to enable the manipulation of a wider variety of objects, as well as integrate active perception and navigational abilities to make our framework more generalizable.
Appendix
@@ Appendix A Sources and Processing of Motion Data
We collected a total of four types of basic reference motion data, including 9 motions related to walking, 5 related to picking up, 4 related to carrying, and 5 related to putting down. All these data are in SMPL format and recorded at 30 fps over 139 frames. They all originate from the ACCAD subset of the AMASS [17] dataset. Additionally, to ensure the stability of cooperative tasks involving multiple individuals, we included data for sidewalk and reverse carry motions. The sidewalk data comes from the CMU subset within AMASS, while reverse carry data was scarce. Therefore, we created reverse carry data by reversing the process of the carry data. In total, we used 26 motion data as references. Additionally, we performed a simple visualization of the extended objects as in Figure 6, which sampled from dataset [6].
@@ B.1 Task Observation
While the root height is measured in the global reference frame, all other components are defined in the frame local to the character. Rotations follow a 6-dimensional representation for continuity [46]. The simulated character aligns with [25, 24, 7, 21], featuring 12 internally movable joints and a total of 28 degrees of freedom.
@@ B.2 Reward Functions
Follow the formulation of the AMP framework [25], the style reward $r^{S}$ is calculated according to the discriminator:
@@ Appendix C Implementation Details
Adopting the methodology of AMP [25], we develop a low-level controller encompassing both policy and discriminator networks. The policy network is bifurcated into a critic and an actor-network, each initiating with a CNN layer and proceeding to two MLP layers configured with [1024, 1024, 512] units. The discriminator network is similarly structured, featuring two MLP layers with [1024, 1024, 512] units. We select PPO [28] as the primary reinforcement learning algorithm, coupled with the Adam optimizer [12] at a learning rate of 2e-5. The only difference between the multi-agent setting and the single-agent setting during training is whether a pre-trained weight is loaded. Our experiments are conducted on the IsaacGym simulator [18] using a single Nvidia GTX 3090Ti GPU. We run 4096 parallel environments across 15,000 epochs, which takes approximately 15 hours to complete.
@@ C.2 Hyperparameters
Following previous work[25, 7, 21], we use the Isaac Gym simulator [18]. The simulation runs at 60Hz and the control policy runs at 30Hz.
@@
Appendix D More Ablation Studies and Visulizations.
Failure case visualization.
@@
Ablation on performance of CooHOI under noisy scenarios.
In CooHOI, all state information is provided using ground truth data from simulators. However, in real-world settings, input data is often noisy and prone to errors. To evaluate the robustness of our framework in such conditions, we introduce random noise into the observation space of our policy and assess its performance under observation noise, as shown in Table 4.
Table 4: Results of our policy under noisy conditions: We tested both single-agent and two-agent box-carrying scenarios. The noise level is defined by the standard deviation of the Gaussian noise used. SR stands for success rate. The definitions of success rate and precision are consistent with those in Section 4.1 of our paper.
Agent # Weight(kg) Noise SR (%) Precision (cm)
1 10 0 96.85 5.76
1 10 1 95.80 6.82
1 10 2 78.56 9.28
1 10 3 60.03 10.75
1 10 4 48.48 8.62
2 20 0 90.33 8.80
2 20 1 90.23 8.96
2 20 2 87.98 8.92
2 20 3 84.86 9.01
2 20 4 79.93 9.28
Training reward curves.
To enhance visualization for ablation studies, we plot the training curves for both the carry reward and held reward in the two-agent training setup. The results, shown in Fig 8, illustrate the effectiveness of our CooHOI framework.